# Orthogonal projectors of binary LCD codes

### v2

@@ 4 Applications
In this section, we illustrate examples where results from the preceding section can be applied.
4.1 Binary codes from the Paley graph
Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. They observed that for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight. Based on this result, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]). The classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights.
On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet. However, it is not difficult to show the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$. One way to see this is to use the Griesmer bound, which tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$. By Lemma 3.6, the minimum weight must be even, and hence the result follows. According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so it is obvious that the code from the Paley graph of order $41$ attains the largest minimum weight. To best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
Remarkably, the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7. We state this fact as the following corollary:
@@ 4.1 Binary codes from the Paley graph
According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism. We constructed binary even LCD codes from the $7152$ graphs, and verified the following:
  • There are mutually inequivalent,
  • Only the code that comes from the Paley graph has the largest minimum weight $10$,
  • All codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$.
By Proposition 3.10, any code from $srg(41,20,9,10)$ must have the minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.7, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
4.2 Optimal LCD codes from regular graphs
In this subsection, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors. For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021]. Let $d(n,k)$ denote the largest minimum weight among all LCD $[n,k]$ code.
Here we consider a construction by circulant matrices. A circulant matrix $C$ of order $n$ is a matrix of the following form
$C=\begin{pmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-2}&c_{n-1}\\ c_{n-1}&c_{0}&c_{1}&\cdots&c_{n-3}&c_{n-2}\\ c_{n-2}&c_{n-1}&c_{0}&\cdots&c_{n-4}&c_{n-3}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ c_{2}&c_{3}&c_{4}&\cdots&c_{0}&c_{1}\\ c_{1}&c_{2}&c_{3}&\cdots&c_{n-1}&c_{0}\\ \end{pmatrix}.$
Circulant matrices are determined by their first rows, and hence there are $2^{n}$ circulant matrices over $\mathbb{F}_{2}$ of order $n$. For $26\leq n\leq 37$, we constructed all circulant matrices over $\mathbb{F}_{2}$ of order $n$, which must be checked further for the condition given in Theorem 3.7. Then by Lemma 3.1 we constructed the binary LCD codes from the orthogonal projectors, and computed their minimum weights. Consequently, we obtain the following theorem:
@@ 4.2 Optimal LCD codes from regular graphs
There are optimal binary LCD $[n,k,d]$ codes for
$(n,k,d)=$ $(31,10,10),(31,11,10),(31,20,6),(33,10,12),$
$(33,11,11),(33,12,10),(33,13,10),(33,20,6),(33,22,6).$
For each of these codes, the first rows of their orthogonal projectors are listed in Table 1. For a binary LCD $[n,k,d]$ code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code. Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2. Consequently, we obtain the following theorem:
Theorem 4.3.
There are optimal binary LCD $[n,k,d]$ codes for
$(n,k,d)=(33,21,6),(34,13,10),(34,22,6).$
Any orthogonal projector mentioned above has constant diagonal entries, and hence any binary LCD code in Theorem 4.2 is either an even code or its dual code. Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction. It is remarkable that adjacency matrices of regular graphs can give LCD codes with the largest minimum weights. Furthermore, we remark that any code in Theorem 4.2 has odd length. This follows by the fact that the dimension of any binary even LCD code is even, and that the dual code of any even code must contain the all-ones vector; if a binary even LCD $[n,k]$ code $C$ has even length, then the all-ones vector must be in $C\cap C^{\perp}$, a contradiction.
Table 1: Optimal binary even LCD codes
$n$ $k$ $d$ the first row
$31$ $10$ $10$ $(0000010001110100001011100010000)$
$31$ $11$ $10$ $(1001011101111110011111101110100)$
$31$ $20$ $6$ $(0001011101111110011111101110100)$
$33$ $10$ $12$ $(000101110111111100111111101110100)$
$33$ $11$ $11$ $(100001010010011000011001001010000)$
$33$ $12$ $10$ $(000101110110111100111101101110100)$
$33$ $13$ $10$ $(100001010011011000011011001010000)$
$33$ $20$ $6$ $(000001010011011000011011001010000)$
$33$ $22$ $6$ $(000001010010011000011001001010000)$
@@ 5 Concluding remarks
In this paper, we proved that binary even LCD codes and some graphs are in one-to-one correspondence in a certain way. Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs. Consequently, we gave an explanation to the observation by Haemers, Peeters, Rijckevorsel [Haemers1999-no]. Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4. We hope that studying such graphs leads to a discovery of a further relation between graphs and even LCD codes.
@@ 6 Acknowledgement
The author would like to thank Stefka Bouyuklieva for several questions and comments that improved the manuscript.

### v3

@@ 5 Applications
Haemers, Peeters, and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. With the help of computers, they observed that
  1. (i)
    for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes;
  2. (ii)
    among such codes the code from the Paley graph had the largest minimum weight.
Based on these results, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]). In this section, we (i) prove that the former statement is true for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$, and (ii) verify the latter statement for $srg(41,20,9,10)$. Note that the result by Brouwer and Eijl [Brouwer1992-nc] shows that for strongly regular graphs $srg(25,12,5,6)$ and $srg(41,20,9,10)$, corresponding binary codes have dimension $12$ and $20$, respectively.
First, we verify that the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$. This can be done easily using the main results of this paper.
@@
Proof.
Let $A$ be an adjacency matrix of $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$. By Theorem 2.4 we have $A^{2}=\frac{q-1}{2}I+\frac{q-5}{4}A+\frac{q-1}{4}(J-I-A)$, where $I$ and $J$ denote the identity matrix and the all-ones matrix of suitable order, respectively. Since we assume that $q\equiv 1\pmod{8}$, $A^{2}=A$ holds when we regard $A$ as a matrix over $\mathbb{F}_{2}$. This fact, together with Theorem 3.6, implies that the resulting code is an even LCD code. Hence the result follows. ∎
Next, we study the second part of the observation above. We remark that the classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights. On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet, making it difficult to prove the problem in a similar way.
@@
The Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$.
Proof.
The result by Brouwer and Eijl [Brouwer1992-nc] shows that for $srg(41,20,9,10)$ any corresponding binary code has dimension $20$, respectively, and the Griesmer bound tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$. Recall that by the proof of Corollary 5.1, any $srg(41,20,9,10)$ gives an even LCD code, and hence the minimum weight must be at most $10$. A computer search verifies that the Paley graph gives a code which attains this bound. This concludes the proof. ∎
According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so we can simply use this fact to prove Theorem 5.2. However, the proof of this bound depends on a lengthy computer search, making it difficult to understand. From this perspective, the proof presented in this paper is more sophisticated. To the best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
Remark 5.3.
It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound. The author thinks that this problem is interesting.
According to [Maksimovic2022-xa], there are at least $7152$ strongly regular graphs with parameters $(41,20,9,10)$ up to isomorphism. We constructed binary even LCD codes from the $7152$ graphs, and verified that (i) there are mutually inequivalent; (ii) only the code that comes from the Paley graph has the largest minimum weight $10$; and (iii) all codes have minimum weight greater than $4$, and the code comes from the following adjacency matrix has the minimum weight $4$. By Proposition 4.2, any code from $srg(41,20,9,10)$ must have minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.6, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
@@ 6 Concluding remarks
In this paper, we have established one-to-one correspondences between binary even LCD codes and certain simple graphs; and ternary LCD codes and certain two-graphs. Based on these results, we further studied the relation between LCD codes and graphs. Our motivation is to study whether a good graph gives a rise to a good code. We verified that for some strongly regular graphs the answer is positive, for (i) the automorphism groups of graphs equal those of the resulting codes, and (ii) the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all codes from $srg(41,20,9,10)$. It is unknown whether there is any other $srg(41,20,9,10)$ than the Paley graph which attains the bound, and the author thinks that this problem is interesting.
@@ 7 Acknowledgement
The author would like to thank Stefka Bouyuklieva for her insightful questions and comments, which improved the manuscript.